OpenStudy (anonymous):

change each logarithmic expression to an exponential expression. ln 5=x

OpenStudy (agreene):

e^5 = e^x

OpenStudy (hoblos):

=> ln 5 = xlne => 5= e^x

OpenStudy (anonymous):

e^x = 5

OpenStudy (anonymous):

\[e^x = 5\]